### v1

# Machine Learning Assisted Adjustment Boosts Inferential Efficiency of Randomized Controlled Trials

Abstract
In this work, we proposed a novel inferential procedure assisted by machine learning based adjustment for randomized control trials. The method was developed under the Rosenbaum’s framework of exact tests in randomized experiments with covariate adjustments. Through extensive simulation experiments, we showed the proposed method can robustly control the type I error and can boost the inference efficiency for a randomized controlled trial (RCT). This advantage was further demonstrated in a real world example. The simplicity and robustness of the proposed method makes it a competitive candidate as a routine inference procedure for RCTs, especially when the number of baseline covariates is large, and when nonlinear association or interaction among covariates is expected. Its application may remarkably reduce the required sample size and cost of RCTs, such as phase III clinical trials.
1 Introduction
The randomized controlled trial (RCT) is the gold standard in determining treatment efficacy. Typically encompassing a substantial cohort size, particularly evident in phase III clinical trials, RCTs are characterized by their extensive participant involvement. However, this rigorous process comes at a considerable cost. Especially for trials on immunotherapy, the cost of some the newest therapies can reach \$850,000 per patient when including the value of medical support necessary to deliver these treatments [1]. Given the significant investment required, maximizing the statistical power of RCTs becomes critical.
A common method used for baseline covariate adjustment is analysis of covariance (ANCOVA), which is essentially a linear model. In 2002, Rosenbaum proposed method of exact permutation inference which is free of distributional assumptions [2]. Although the method discussed was based on ordinary least square (OLS) linear regression, the author raised possibility of using more general forms of covariate adjustment. Nowadays, due to the ever decreasing cost of medical testing and sequencing techniques, the number of baseline data available in clinical trials has been growing dramatically. The rich data can provide a much finer picture of individuals and thus capture larger variation in the outcomes. This makes it possible to design inferential procedures with superior efficiency. However, the traditional adjustment methods using OLS method quickly break down when the number of covariates increases to a relatively large scale. Even if penalized regressions [3, 4, 5, 6] can acommodate high-dimensional data, such linear models are neither capable of handling nonlinear associations nor complex interactions.
On the other hand, machine learning techniques have been successfully applied in numerous settings during the last decade. Supervised learning methods such as random forests (RFs), boosting machines, neural networks [7, 5, 8, 11] are known to be capable of handling high-dimensional data and model complex functional forms. In this work, we find the novel application of RF under the Rosenbaum’s framework is a valid and highly efficient adjustment method to boost the efficiency of inference for a RCT. The simplicity and robustness of the method makes it a potential routine inferential procedure for RCTs, especially when the number of baseline measurements is large, and when nonlinear association or between-covariate interaction is expected. The RF is used as a proof of concept, but it can be extended to more general modeling approches. Its application may remarkably reduce the required sample size and cost of RCTs, especially for phase III clinical trials.
2 Methods
In this work, we focus on the RCT with continuous outcomes and objective to compare two group means. Our method was based on the Rosenbaum’s framework. For details, please see Rosenbaum [2], and we here give a brief review. Suppose there are $n$ subjects, $j=1,\dots,n$, and the response of subject $j$ would be $r_{T_{j}}$ or $r_{C_{j}}$ if $j$ were assigned to treatment or control, respectively [9, 10]. The treatment effect can be written as $r_{T_{j}}-r_{C_{j}}$, and it can never be observed because subject $j$ can only receive one of the treatments. The treatment effect is additive if the treatment causes the response to change by a fixed amount $\tau$ for every $j$. The covariate vector $x_{j}$ describes the baseline characteristic of $j$. The variable $Z_{j}$ is the treatment indicator, where $Z_{j}=1,0$ indicates $j$ is assigned to the treatment or control group, respectively. Thus, the response of $j$ can be written as $R_{j}=Z_{j}r_{T_{j}}+(1-Z_{j})r_{C_{j}}$. Following Rosenbaum [2], we denote $\mathbf{Z}=(Z_{1},Z_{2},\dots,Z_{n})^{T}$, $\mathbf{R}=(R_{1},R_{2},\dots,R_{n})^{T}$, $\mathbf{r}_{c}=(r_{C_{1}},r_{C_{2}},\dots,r_{C_{n}})^{T}$ and use $\mathbf{X}$ as the covariate matrix whose $j$th row is $\mathbf{x}_{j}^{T}$. The fixed but unobserved vector $\mathbf{r}_{C}$ are potential outcomes of all subjects if assigned to the controlled group, which and can be expressed as $\mathbf{r}_{C}=\mathbf{R}-\tau\mathbf{Z}$. Under this framework, the only stochastic component is the random assignment $\mathbf{Z}$. The test of $H_{0}:\tau=\tau_{0}$ can be performed using the statistic $t(\mathbf{Z},\mathbf{R}-\tau_{0}\mathbf{Z})$, where $\mathbf{R}-\tau_{0}\mathbf{Z}$ equals $\mathbf{r}_{C}$ under $H_{0}$ and additive treatment effects. Under this framework, the exact inference is the randomization inference derived from the randomization distribution of the statistics. The parametric distributions such as the normal distribution are used as approximations to randomization distributions [2].
When covariate information $\mathbf{X}$ is available and correlate with the response variable $\mathbf{r}_{C}$, then we can write $\mathbf{r}_{C}=g(\mathbf{X})+\mathbf{e}$. The residual $\mathbf{e}$ is expected to have less variation than $\mathbf{r}_{c}$ if the baseline information explain part of the variation in the outcome. Note that if $H_{0}$ is true, we will have $\mathbf{r}_{C}=\mathbf{R}-\tau_{0}\mathbf{Z}$, and the residuals $\mathbf{e}$ can be used in place of $\mathbf{R}-\tau_{0}\mathbf{Z}$ for inference. Namely, the test statistic can be modified to $t(\mathbf{Z},\mathbf{e})$, which can lead to an improvement in efficiency due to smaller variation in $\mathbf{e}$.
Linear models are natural choice of $f$, but they tend to fail when the association is complex or the data is high dimensional. Here we investigate whether using modern machine learning method will achieve better efficiency for testing $H_{0}:\tau=\tau_{0}$ as compared to the straight two group comparison, while controlling the type I error rate properly. In this work, we select RF as the modeling method. For simplicity, let $\mathbf{Y}=(Y_{1},Y_{2},\dots Y_{3})^{T}$ denote the vector of continuous outcomes and note that we are interested in a regression task. The RF model is trained by repeatedly drawing bootstrap sample from the training data $(\mathbf{X},\mathbf{Y})$ for $B$ times. For the $b$th bootstrapped data $(\mathbf{X},\mathbf{Y})^{*}_{b}$ grow a random forest tree $T_{b}$ through recursive partitioning. A key feature of RF is that only $m<p$ variables will be randomly selected as candidates for each splitting, which reduces the correlations between trees and thereby improves the benefits of variance reduction by averaging. For a new observation $x$, the predicted response will be $\hat{f}(x)=\frac{1}{B}\sum_{b=1}^{B}T_{b}(x)$. For regression, the default $m=\lfloor\sqrt{p}\rfloor$, and the minimum node size is 5 [8].
The RF approach was selected as a proof of concept due to its minimal requirement of parameter tuning and invariance to monotonic transformation of features [7]. More importantly, the out-of-bag (OOB) predictions can be used to attenuate the overfitting, thus avoiding manually splitting the data into training and validation set. Such simplicity makes it ideal for RCTs which typically requires detailed preplanning.
The proposed procedure for testing $H_{0}:\tau=\tau_{0}$ is straight forward:
  1. 1.
    Use the data of all subjects and build a RF model using $\mathbf{X}$ as the features and $\mathbf{R-\tau_{0}\mathbf{Z}}$ as the response $\mathbf{Y}$.
  2. 2.
    Obtain the OOB predictions for all subjects and calculate the residuals $\mathbf{e}$.
  3. 3.
    Perform the permutation test, Wilcoxon rank-sum test or two sample $t$-test based on $\mathbf{e}$ and $\mathbf{Z}$.
  4. 4.
    Estimate the treatment effect, standard error or confidence interval by treating $\mathbf{e}$ as the outcome variable.
Under randomization and the assumption that the nonlinear models shift by some fixed amount $\bar{e}_{1}-\bar{e}_{0}=\bar{R}_{1}-\bar{R}_{0}$, so the average difference of the residuals by treatment group is the estimated treatment effect, while the residuals have a less dispersion. In machine learning, overfitting is a commonly encountered problem when a model has excessive flexibility and fits closely to the training data set. Such models are characterized with high variance and generalize poorly to new observations. This can also occur in modeling the relation between $\mathbf{X}$ and $\mathbf{R}-\tau_{0}\mathbf{Z}$ and lead to a biased estimation of the treatment effect.
3 Simulations
To investigate the performance of the proposed method, we conducted simulation studies under various scenarios for testing a one-sided hypothesis $H_{0}:\tau=0$ versus $H_{1}:\tau>0$. The results can be readily extended to two-sided tests. In each scenario, we generated a sample of $N=50,100,200$ and $400$ independent subjects, who are randomized into balanced two groups. There are $p=40$ mutually independent covariates following standard normal distributions. In each setting 10,000 Monte Carlo simulations were performed. The outcome $y$ was generated using the four different models:
  1. 1.
    Model 1 (primarily nonlinear no interaction):
    $y=\tau Z+\beta\sigma(x_{1}/2)+\beta x_{2}^{2}+\beta cos(x_{3})+\beta x_{4}+\epsilon,$
  2. 2.
    Model 2 (primarily nonlinear with interaction):
    $y=\tau Z+\beta\sigma(x_{1}/2)+\beta x_{2}^{2}+\beta cos(x_{3})+\beta sign[cos(x_{3})]x_{4}+\epsilon,$
  3. 3.
    Model 3 (primarily linear no interaction):
    $y=\tau Z+\beta\sigma(x_{1}/2)+\beta x_{2}+\beta x_{3}+\beta x_{4}+\epsilon,$
  4. 4.
    Model 4 (primarily nonlinear with by treatment interaction):
    $y=\tau Z+\beta[\sigma(x_{1}/2)-0.5]Z+\beta x_{2}+\beta x_{3}+\beta x_{4}+\epsilon,$
where $Z$ is the treatment indicator, $\tau$ is the treatment effect, $\sigma(x)=\exp(x)(1+\exp(x))^{-1}$ is the sigmoid function, $x_{k}$ is the $k$th covariate, and $\epsilon$ is the error term. The subscript for subject is omitted. We examined $\beta=0.2,0.5,0.8$ to represent different association strength between covariates and the outcome, as well as error terms from standard normal, log-normal, and Gumbel distributions. The straight two-sample $t$-test and Wilcoxon rank-sum test are commonly used analytical methods. We also included $t$-test and Wilcoxon rank-sum test on the residuals from linear regression on all baseline covariates ($t$-test-LM and Wilcoxon-LM). Finally, we included the proposed RF adjusted methods, which use the $t$-test and Wilcoxon rank-rum test on the residuals from RF model ($t$-test-RF and Wilcoxon-RF). We examined the type I error control with $\tau=0$, and the statistical power with $\tau=0.3$ and $0.6$.
Figures 1$\sim$4 show the simulation results from models with normal error terms. It is shows that all tests robustly control the type I error rates when $\tau=0$ in all settings. In models with primary nonlinear associations (Model 1, 2, 4), the RF adjusted tests ($t$-test-RF and Wilcoxon-RF) show uniformly higher power than the other models. In Figure 1, bottom right panel, the RF tests attains over 90% power at around $N=200$, while the other test would require at least 100 more subjects to achieve the same power. In models with primary linear associations (Figure 3, Model 3), the performance of RF tests is mostly comparable with that of the linear model adjusted models. Only when the association is very strong (Beta=0.8), the linear model methods will have a advantage of around 7%. It should be noted that under some scenarios the linear model tend to hurt the efficiency, making its power lower than the straight $t$-test and Wilcoxon test, even when the true association is primarily linear. On the other hand, the RF tests consistently have similar or better performance versus the unadjusted tests. This suggests the application of RF adjustment is generally safe and unlikely to result in a loss of efficiency. Figure 5 shows that when the distribution of error is non-normal, all tests still control the type I error robustly. Meanwhile, the RF tests still exhibit similar advantage over other methods, but the Wilcoxon-RF shows a further improvement over the $t$-test-RF under non-normality. The results on cases with log-normal errors are similar (Supplement). In addition, although the Rosenbaum’s framework assumes an additive treatment effect, the proposed method show equally well type I error control and efficiency improvement when there are interactions between treatment and covariates (Figure 4). The results are similar is the randomization is unbalanced (data not shown). Overall, our simulations suggest Wilcoxon-RF test as the best option among the methods tested.
Figure 1: Type I error (Tau=0) and statistical power of the testing procedures under different settings of Model 1 and with error terms from standard normal distributions.
Figure 2: Type I error (Tau=0) and statistical power of the testing procedures under different settings of Model 2 and with error terms from standard normal distributions.
Figure 3: Type I error (Tau=0) and statistical power of the testing procedures under different settings of Model 3 and with error terms from standard normal distributions.
Figure 4: Type I error (Tau=0) and statistical power of the testing procedures under different settings of Model 4 and with error terms from standard normal distributions.
Figure 5: Type I error (Tau=0) and statistical power of the testing procedures under different settings of Model 4 and with error terms from standard Gumbel distributions.
4 Diet Intervention Data Example
In this example, the investigators studied whether a brief diet intervention can reduce the symptoms of depression in young adults. There were 76 particpants randomly allocated to a brief 3-week diet intervention (Diet Group) or a habitual diet control group (Control Group), with $n=38$ per group. The primary and secondary outcome measures were assessed at both baseline and after the intervention, which included symptoms of depression (Centre for Epidemiological Studies Depression Scale, CESD-R; and Depression Anxiety and Stress Scale-21 depression subscale, DASS-21-D), current mood (Profile of Mood States, POMS), self-efficacy (New General Self-Efficacy Scale, GSES) and memory (Hopkins Verbal Learning Test). There are a total of 23 baseline variables can be used as covariates. The GSES score were log-transformed as specified in the article. The ANCOVA was used for the analysis of both primary and secondary outcomes with the baseline scores adjusted as covariates. The results showed that the Diet group had significantly lower self-reported depression symptoms than the Control Group on the CESD-R and DASS-21 depression subscale. However, none of the POMS scores showed significant difference. Here we used our method (Wilcoxon-RF) to re-analyzed the difference of POMS fatigue score and two-sided tests were used. The results in Table 1 show that only the Wilcoxon-RF suggest there is a significant difference in the post-intervention fatigue scores between two groups. This demonstrates the higher statistical power of the proposed method. Based on the residuals, the final estimated treatment effect is -1.57, with 95% confidence interval of (-3.01, -0.05). The variable importance from RF suggests baseline depression is the most important predictor. The partial dependence plot (Figure 6) suggests the post-intervention fatigue does not change with the baseline depression within the range of $0\sim 5$, but starts to show a linear trend in the range of $6\sim 12$, and then plateaus. Such sigmoid shape may not be well described by a linear model, but can be better captured by a nonlinear model like RF, which explains why our method achieves better efficiency.
Method $p$-value
$t$-test 0.2907
Wilcoxon 0.2452
$t$-test-LM 0.3112
Wilcoxon-LM 0.3661
$t$-test-RF 0.1248
Wilcoxon-RF 0.0451
Table 1: Results of testing methods applied to the diet intervention data for comparing post-intervention POMS fatigue score between the Diet and Control groups. All tests are two-sided.
Figure 6: The partial dependence of Day 21 fatigue score and baseline depression score. Blue dashed line is the smoothed partial dependence by Loess smoother.
5 Discussion
In this work, we presented a novel inferential procedure for RCTs assisted by an machine learning adjustment. Through a set of simulation experiments and a real world example, we showed the proposed method remarkably boosts the statistical power and can reduce 1/3 sample size requirement under certain scenarios. Such improvement is grounded in the capability of machine learning in capturing nonlinear relationships and higher order interactions, which cannot be fully exploited by linear models. We also demonstrated that the proposed method is generally safe to use in practice because the type I error is robustly controlled. Meanwhile, under the scenarios tested, it never shows worse performance than the straight two-sample tests without adjustments. In contrary, the methods using adjustment by linear regression can sometimes harm the efficiency, which can be attributed to its instability when $p$ is large relatively to $N$. Shrinkage methods such as Ridge [3], Lasso [4] or Elastic net [6] may be used, but they all require careful tuning of hyperparameters, and they do not address the difficulty brought by non-linearities and interactions.
The RF is often considered an “off-the-shelf” supervised learning method due to its simplicity of usage. Typically with the recommended settings one can obtain a model with good performance [8]. This makes it a good choice for RCTs which typically requires detailed analysis plans to be pre-specified in the study protocols. The usage of OOB observations further makes explicit training, validation set splitting unnecessary. Therefore, in an ideal scenario, the inference procedure can be completed in a single run of RF model training. Another motivation for using RF is that it usually shows better or comparable prediction performance than neural networks in the range of sample sizes typically used in a randomized study. Deep learning models typically require larger sample size to gain advantage, and it usually requires extensive tuning [11], which can complicate the inference procedure and is hard to plan beforehand. However, no model can be optimal under all scenarios. Although RF is used as a proof of concept in our study, other modeling methods can be superior to RF in many settings. Therefore, investigation of the proposed approach to incorporate other machine learning techniques is important.
Another practical challenge is how to translate the gain in efficiency into reduction of sample size. This requires the knowledge of the variation that can be explained by the machine learning model at the planning stage. In medical research, a phase II trial is conducted to preliminarily assess the efficacy before conducting a large scale phase III RCT. The data collected from the phase II trial provides important parameters that can guide the design of the phase III trial. One can simply use the two-sample $t$-test formula as an approximation for the exact test, and estimate the proportion of variance that will be reduced by including the covariates based on assumptions or data available, e.g. from a phase II trial. Approximately, if variance explained by the model is $\gamma$, then one would expect a reduction of $N\gamma$ in sample size requirement, where $N$ is the sample size needed using a unadjusted two-sample test. The parameters may also be used to run simulations to estiamte the sample sizes required for the machine learning adjusted tests. Another option is an interim analysis on the efficiency gain and adjust the sample size accordingly. Such strategy may require further methodological investigation.